An infinitely long line of charge carries 75 pC along each 1 meter of length. Find the electric field 0.40 m from the line of charge.

1
Expert's answer
2022-04-13T09:20:55-0400
"E=\frac{\lambda}{2\pi\epsilon_0 r}"

"E=\frac{75*10^{-12}}{2\pi*8.85*10^{-12}*0.40}=3.4\:\rm N/C"